### v1

@@ 3 The GRAND detection unit
The GRAND detection unit (DU) consists of an antenna—the HorizonAntenna—and the electronics to process, store, and communicate the signals detected. The DU design is optimized to detect radio signals that reach the array from near-horizontal directions, as expected from inclined air showers initiated by cosmic rays and neutrinos [6, 7].
Although the DU design is common to GP300 and G@A, its built-in flexibility allows it to be modified and adapted, a feature exploited in the instrumentation of the three GRAND prototype arrays. During their deployment and commissioning, different configurations of the hardware and software were installed at each array for two reasons. First, adapting the instrumentation was necessary for the setups to operate under the different environmental conditions present at each site. Site-specific adaptations include heat and moisture countermeasures, hibernation conditions, and tailored trigger settings. Second, exploring alternative configurations of the instrumentation allowed us to assess the performance of alternative hardware and software choices, and thus field-test options for later stages of GRAND.
@@ 3.1 Mechanical structure
Figure 3 shows the mechanical structure of the DUs deployed at GP300. A 3.5-m pole holds a “nut” onto which the five arms of an antenna are mounted. The nut houses the low- noise amplifiers (LNAs). In total, three cables run from the LNAs through the main pole to the GRAND front-end board (FEB), which is kept inside a box within a ventilation casing on the back of the triangular-shaped base on the ground. A mesh communications antenna serving WiFi communication with the corresponding central station is mounted about halfway on the main pole. The power for all elements of the DU is provided by a solar panel, connected through a charge controller to a battery. The solar panel functions as a door that allows access to both. In addition, the pole can be lowered to facilitate access to the nut.
Figure 4 shows the mechanical structure of the DUs deployed at G@A, closely resembling those at GP300. The original AERA units were modified to better resemble the GRAND DU design. In particular, the smaller vertical pole was extended by fitting a hollow aluminum tube inside the existing pole, thus raising the antenna to a height of 3 m—50 cm lower than the antennas at GP300. The original AERA solar panel and the metal enclosure were kept, with the GRAND FEB box, battery, and charge controller stored inside. In the following sections, we discuss the characteristics of the antenna, LNAs, the GRAND FEB, power supply, and communications setup.
@@ 3.2 Antenna and low-noise amplifier
Antenna.—Figure 5, left, shows that the antenna itself consists of five arms: two dipoles in the North-South and East-West directions, and a monopole in the vertical direction. The dipole size is roughly 1.4 m, to match the GRAND working frequency band of 50–200 MHz. Figure 5, right, shows the acrylic nut with an internal connector to which the antenna arms are attached. The LNA board is housed in the metal base of the nut and is powered by a coaxial cable that also transports received signals to the FEB.
@@
Figure 5: Schematic diagrams of the GRAND antenna and nut. Left: The antenna, with three arms: two dipoles in the North-South and East-West directions, and a monopole in the vertical one. Right: The nut: the acrylic, in blue, holds the antenna arms and protects the interior metal connectors, inside the support, in yellow. These connect the antenna arms to the LNA board, held at the metal base. See Section 3.2 for details.
@@
Figure 6: Simulated vector effective length (VEL) of the GRAND HorizonAntenna. The magnitude of the VEL, $|L_{\rm eff}|$ as a function of frequency is different for each of the three antenna arms: North-South (left), East-West (center), and vertical (right). In this figure, as an example, we show results for an azimuth angle of $45^{\circ}$ measured from the North, and for three illustrative choices of the zenith angle: $60^{\circ}$, $75^{\circ}$, and $85^{\circ}$. See Section 3.2 for details.
Figure 7: Simulated group delay of the GRAND HorizonAntenna. The group delay as a function of frequency is shown for each of the three antenna arms, as in Fig. 6, and for the same illustrative choices of the azimuth and zenith angle. See Section 3.2 for details.
@@
Figure 6 shows the simulated response of the GRAND HorizonAntenna to an incoming radio signal, expressed in the form of the magnitude of the vector effective length (VEL), $\mathbf{L}_{\rm eff}$. The VEL is the transfer function between the incoming electric field, $\mathbf{E}$, and the open-circuit voltage at the antenna terminals, $\mathbf{V}_{\rm oc}$, i.e., $\mathbf{V}_{\rm oc}=\mathbf{L}_{\rm eff}\cdot\mathbf{E}$. The larger the magnitude of the VEL, the higher the voltage generated by the antenna for a given incident electric field, provided that the electric field polarization and the VEL are aligned; see, e.g., Eq. (1) in Ref. [48].
@@
For the HorizonAntenna, the VEL was simulated using the Ansys HFSS (high-frequency structure simulator) program, with a setup that included the complete mechanical structure of the DU. The resulting VEL is an elaborate function of frequency, zenith, and azimuth angle that fully characterizes the response of the antenna to an arbitrary incoming plane-wave signal; see Fig. 5 in Ref. [48]. This is an essential ingredient to produce simulated measurements and interpret real ones. For the HorizonAntenna, Fig. 6 shows that the North-South and East-West arms have similar VELs, since the dipoles are of equal size, while the vertical arm has a markedly different VEL, likely due to the different size of the monopole, its different orientation, and the reflection of radio waves off the ground. Given the complexity of the VEL of the HorizonAntenna, its validation in the field remains to be completed; a calibration campaign is planned to verify our simulations.
Figure 8: Simulated linear gain of the GRAND HorizonAntenna. The linear gain as a function of frequency is shown for each of the three antenna arms, as in Figs. 6 and 7, for the same illustrative choices of zenith angle, and for a fixed azimuth angle of $45^{\circ}$. See Section 3.2 for details.
@@
Figure 7 shows the group delay of the HorizonAntenna, i.e., the derivative with respect to the frequency of the phase shift of the transfer function of the DU (Section 3.4). A varying group delay throughout the working frequency range can lead to signal dispersion and loss of amplitude due to interference effects. For air-shower detection and reconstruction, the shape and amplitude of the signal are important assets used, for example, in triggering algorithms. For the HorizonAntenna, the group delay is nearly flat in frequency for the three antenna arms, with differences well below 20 ns that do not significantly distort the expected pulse shape. Given the frequency bandwidth of 200 MHz, group delay variations smaller than its inverse (5 ns) are acceptable. While there are sharp peaks in the group delay, they are narrow enough to have only a small effect on signal power. The antenna response is embedded in the GRANDlib processing pipeline [48] (Section 3.4) and is used in the reconstruction of the properties of showers detected by GRAND.
@@
Figure 9: Low-noise amplifier (LNA) board deployed at GRANDProto300 (left) and GRAND@Auger (right). The main components of the boards are: (1) connectors to the antenna, (2) balun, (3) matching network, (4) amplifier, and (5) power. Although the components are in similar locations, their specifications are not the same in both boards. See Section 3.2 for details.
@@
LNA.—Figure 9 shows the two designs of the LNA boards deployed in GP300 and G@A. For GP300, the LNA board was designed to provide a stable gain of about 20 dB in the range of 50–200 MHz. As shown in Fig. 9, left, in the GP300 board the X and Y paths are connected to the dipoles and the Z path, to the monopole. When powered with 5 V from the FEB (more on it later), the LNA board draws 47 mA for each channel, leading to a power consumption of about 235 mW for the full board. An electrostatic discharge protection structure guards the LNA from damage due to static electricity. The balun converts differential signals into single-ended signals and performs impedance transformation. The matching network, located at the front-end of the LNA, converts the impedance to 50 $\Omega$, minimizing impedance matching issues at the LNA input port.
@@
For G@A, the LNA design was focused on providing higher gain at a low power consumption. Our design provides a stable gain of about 30 dB for each of the X, Y, Z channels, in the same range as GP300, 50–200 MHz. The noisier radio environment of G@A compared to GP300 leads to additional noise pickup. Were it not for the high-gain LNA, this noise would be amplified by the variable gain amplifier on the FEB. Like in GP300, the board is powered with 5 V by the GRAND FEB; each channel draws 40 mA and the total power consumption of the board is 200 mW. As shown in Fig. 9, right, the path lengths for the X, Y, and Z channels in the LNA board are exactly the same, leading to equal signal time delays for all channels. This allows for the study of the time dependence of the polarization in an unbiased way, avoiding additional software corrections to the timing of the individual polarization directions. Given the weather conditions in Argentina, we implemented moisture countermeasures to prevent water damage: the nuts are sealed with an additional layer of silicone, and the LNA boards are coated with transparent lacquer.
@@
The main differences in the output signals provided by the two designs of the LNA board are the gain and the matching network. For GP300, the LNA has uniform matching across the full frequency range. For G@A model, its matching is optimized for higher frequencies, i.e., above 80 MHz. The difference in maximum gain for the LNA boards is compensated by the amplifier in the GRAND FEB, as described in the next section.
3.3 Front-end board and firmware
@@
Figure 10: The GRAND front-end board (FEB). The main components: (1) signal input and filter chain, with four inputs available; (2) analog-to-digital (ADC) chip; (3) system on chip: field-programmable gate array (FPGA) and central processing unit (CPU) ; (4) Global Positioning System (GPS) chip and connector; (5) Ethernet chip; (6) clock; (7) power supply for the digital part of the board; (8) power supply for the analog part of the board; and (9) power connector and switch. See Section 3.3 for details.
@@
Figure 11: Measured response of the analog filtering on the prototype GRAND FEB with a set 3 dB VGA gain. The plateau at around 3 dB indicates that the signal is properly amplified within the GRAND working frequency range of 50–200 MHz. See Section 3.3 for details.
@@
Figure 10 shows the prototype GRAND FEB used in GP300 and G@A. We describe its primary components we below.
Analog chain.—The analog chain, component (1) in Fig. 10, serves multiple purposes. First, the analog signals, coming from the LNA, are shaped through a fifth-order elliptical 30–200 MHz-bandpass filter. In the following, they pass through the variable gain amplifier (VGA), whose setting is controlled by the digital part of the board. The gain can vary between -10 and +20 dB. In GP300, we use a VGA gain of 20 dB; in G@A, we use a 0 dB gain. In addition, the analog chain supplies power to the front-end LNA boards (Section 3.2) through a bias-T.
Figure 11 shows the measured analog response of the FEB board. For this figure, the VGA gain of the FEB was set at 3 dB before the measurement. The response was then measured to be nearly constant in the GRAND working frequency range of 50–200 MHz, with a value of approximately 3 dB, indicating that the signal is amplified properly.
@@
ADC chip.—The FEBs digitize the signals after they pass the analog chain. The digitization uses a 14-bit 500 MSPS, 4-channel analog-to-digital converter (ADC; Analog Devices, AD9694), component (2) in Fig. 10, in differential mode with an input voltage range of $\pm 0.9$ V. This setting allows us to measure the background noise at around 30 ADC counts, leaving a dynamic range of about 9 effective bits. This yields a compromise between accurate noise determination and maximum signal detection that allows us to use the variations in the Galactic background as a monitoring and calibration tool.
@@
FPGA and CPU chip.—The digitized signals are moved into a system on a chip (SoC) (Xilinx, Zynq Ultrascale+, XCZU7CG-2FBVB900E), component (3) in Fig. 10. It consists of a field-programmable gate array (FPGA) combined with four embedded hardcore central processing units (CPUs), two ARM-A53 and two ARM-R5. The SoC performs tasks such as event triggering, event building, communication, and buffering. These functions are divided into tasks performed by the firmware that runs on the FPGA, and tasks performed by the Petalinux (embedded Linux system) running on the CPU.
The firmware continually retrieves the digital data from the ADC, and can be configured to digitally shape this signal. For instance, currently a total of four filters can be applied to the data of each channel, and dynamic baseline subtraction is available. It is possible to select raw or shaped data for the final readout. If signal shaping is selected, the shaped signal is used for triggering to select interesting events. Otherwise, raw data are used. Triggering is elaborated on in in Section 4.
GPS and slow control parameters.—After triggering, the firmware builds the event. Data from the three input channels are combined with timing information from the GPS chip (Trimble RES SMT360), component (4) in Fig.10, and data from slow control parameters, such as air pressure and battery voltage. These are written in an event header, in which all configuration parameters are also stored. The local software inside the CPU is notified that an event is present and starts the transfer into the local board memory.
Triggering and data transfer.—Two firmware versions were tested, one at GP300 and the other at G@A. They differ in the method of data transfer into memory. The firmware version deployed at GP300 starts by reading event packages by the Advanced eXtensible Interface (AXI) bus, as long as the first-in-first-out queue for storing event data is not empty. Given the quieter radio environment of GP300, the AXI speed is enough to deal with the trigger rate.
However, the higher trigger rate at the G@A site required a faster memory transfer method. Therefore, at G@A, the software uses direct memory access (DMA transfer), moving all parts of each triggered event into double data rate (DDR) memory, which has a total size of 512 MB. The firmware may then apply a secondary “high-level” trigger algorithm to further reject uninteresting events. Afterwards, it saves the starting location of each event in order to easily identify in memory those events that will be required to be saved centrally. Then, both firmware versions send the timing information to the central DAQ software and, upon request by the central DAQ, the complete event can be transferred wirelessly via the Ethernet chip, component (5) in Fig. 10.
@@
The FEB box.—As described in Section 3.1, the FEB is stored in a sealed metal box. The FEB box also acts as a Faraday cage, preventing the electronic noise generated by the FEB components from interfering with the measurements. Due to the high temperatures during the summer time at the GP300 and G@A sites, it was necessary to adapt the FEB boxes to prevent the board from overheating. In GP300, a heat dissipation structure was added to the lid of the box. In G@A, the top lids of the FEB boxes were replaced with a metal mesh. A mesh could not be used in GP300 because sandstorms in the Gobi desert would introduce sand into the electronics. In addition, G@A was fitted with an automatic hibernation switch based on the ADC chip temperature.
3.4 Overall system performance

### v2

@@ 3 The GRAND detection unit
The GRAND detection unit (DU) consists of an antenna—the Horizon Antenna—and the electronics to process, store, and communicate the measured signals. The DU design is optimized to detect radio-frequency signals that reach the array from near-horizontal directions, as expected from inclined air showers initiated by cosmic rays and neutrinos [28, 49].
The GRAND DU design has built-in flexibility, allowing it to be modified and adapted if needed. For instance, this feature was exploited in G@A when the original DU design was altered to accommodate it to the existent mechanical structure of AERA. In addition, the different environmental conditions at each site required specific instrumentation adjustments. Site-specific adaptations include heat and moisture countermeasures, hibernation conditions, and tailored trigger settings. We also profited on this opportunity to assess the performance of alternative hardware and software choices and thus field-test options for later stages of GRAND.
@@ 3.1 Mechanical structure
Figure 3 shows the mechanical structure of the DUs deployed at GP300. A 3.5-m pole holds a “nut” onto which the five arms of an antenna are mounted. The nut houses the Low-Noise Amplifiers (LNAs). In total, three cables run from the LNAs through the main pole to the GRAND Front-End Board (FEB), which is kept inside a box within a ventilation casing on the back of the triangular-shaped base on the ground. A mesh communications antenna serving WiFi communication with the corresponding central station is mounted about halfway on the main pole. The power for all elements of the DU is provided by a solar panel connected through a charge controller to a battery. The solar panel functions as a door that allows access to both. In addition, the pole can be lowered to facilitate access to the nut.
Figure 4 shows the mechanical structure of the DUs deployed in G@A, closely resembling those in GP300, modified to better fit the original AERA stations. In particular, the smaller vertical pole was extended by fitting a hollow aluminum tube inside the existing pole, thus raising the antenna to a height of 3 m—50 cm lower than the antennas at GP300. Raising the antenna 3.5 m above ground may have caused vibrations that the setup could not sustain. The original AERA solar panel and the metal enclosure were kept, with the GRAND FEB box, battery, and charge controller stored inside. In the following sections, we discuss the characteristics of the antenna, the LNAs, the GRAND FEB, power supply, and communications setup.
@@ 3.2 Antenna and Low-Noise Amplifier
Antenna.—Figure 5, left, shows that the GRAND Horizon Antenna consists of five arms: two dipoles in the North-South and East-West directions, and a monopole in the vertical direction. The dipole size is roughly 1.4 m, to match the GRAND working frequency band of 50–200 MHz. Figure 5, right, shows the acrylic nut with an internal connector to which the antenna arms are attached. The LNA board is housed in the metal base of the nut and is powered by a coaxial cable that also transports received signals to the FEB. This structure is common to both prototype arrays.
@@
Figure 5: Schematic diagrams of the GRAND Horizon Antenna and nut. Left: The antenna, with three arms: two dipoles in the North-South and East-West directions, and a monopole in the vertical one. Right: The nut: the acrylic, in blue, holds the antenna arms and protects the interior metal connectors, inside the support, in yellow. These connect the antenna arms to the LNA board, held at the metal base, in green.
@@
Figure 6: Simulated Vector Effective Length (VEL) (top), group delay (middle) and linear gain (bottom) of the GRAND Horizon Antenna. Each property is shown as a function of frequency for the three antenna polarizations: North-South (left), East-West (center), and vertical (right). We show results for an azimuth angle of $45^{\circ}$ measured counter-clockwise from the North, and for three illustrative choices of the zenith angle: $60^{\circ}$, $75^{\circ}$, and $85^{\circ}$.
@@
Figure 6, top, shows the simulated response of the GRAND Horizon Antenna to an incoming radio-frequency signal, expressed in the form of the magnitude of the Vector Effective Length (VEL), $\mathbf{L}_{\rm eff}$. The VEL is the transfer function between the incoming electric field, $\mathbf{E}$, and the open-circuit voltage at the antenna terminals, $V_{\rm oc}$, i.e.,
$V_{\rm oc}=\mathbf{L}_{\rm eff}\cdot\mathbf{E}.$ (3.1)
The larger the magnitude of the VEL, the higher the voltage generated by the antenna for a given incident electric field, provided that the electric field polarization and the VEL are aligned.
@@
For the Horizon Antenna, the VEL was simulated using the Ansys HFSS (High-Frequency Structure Simulator) program, with a setup that included the complete mechanical structure of the DU. The resulting VEL is an elaborate function of frequency, zenith, and azimuth angle that fully characterizes the antenna response to an arbitrary incoming plane-wave signal. This is essential for producing simulated measurements and interpreting real ones. For the Horizon Antenna, Fig. 6, top, shows that the North-South and East-West arms have comparable VELs, expected given the dipoles are of equal size. The vertical arm differs likely due to the different size of the monopole, its different orientation, and the reflection of radio-frequency waves off the ground.
Given the complexity of the VEL of the Horizon Antenna, its validation in the field remains to be completed. A calibration campaign is planned to verify our simulations. It will be performed by sending radio-frequency waves of frequencies between 50-200 MHz, with a spacing of 1 MHz, to the GRAND Horizon Antennas. The emitting antenna will have a known location and will be movable (by truck or drone). In this way, we will have the antenna response dependency on frequency, azimuth, and zenith angles fully characterized. A comparison with the Galactic noise measurement will confirm the absolute value of the calibration up to  10%. This will be an improvement over a simulation-only estimate of about a factor of three.
@@
Figure 6, middle, shows the group delay of the Horizon Antenna, i.e., the derivative with respect to the frequency of the phase shift of the DU transfer function. A varying group delay throughout the working frequency range can lead to signal dispersion and loss of amplitude due to interference effects. For air-shower radio-detection and reconstruction, the shape and amplitude of the signal are important assets used, for example, in triggering algorithms. For the Horizon Antenna, the group delay is nearly flat in frequency for the three antenna polarizations, with differences well below 20 ns that do not significantly distort the expected pulse shape. Given the frequency bandwidth of 200 MHz, group delay variations smaller than its inverse (5 ns) are acceptable. Although there are sharp peaks in the group delay, they are narrow enough to have only a small effect on signal power. The antenna response is embedded in the GRANDlib processing pipeline [13] and is used in the reconstruction of the properties of the air showers detected by GRAND.
@@
Figure 7: Low-Noise Amplifier (LNA) board deployed at GRANDProto300 (left) and GRAND@Auger (right). The main components of the boards are: (1) connectors to the antenna, (2) balun, (3) matching network, (4) amplifier, and (5) power. Although the components are in similar locations, their specifications are not the same in both boards.
@@
LNA.—Figure 7 shows the two designs of the LNA boards deployed in GP300 and G@A. For GP300, the LNA board was designed to provide a stable gain of about 20 dB in the range of 50–200 MHz. As shown in Fig. 7, left, in the GP300 board the X and Y inputs are connected to the dipoles and the Z input, to the monopole. When powered with 5 V from the FEB, the LNA board draws 47 mA for each channel, which leads to a power consumption of about 235 mW for the whole board. An electrostatic discharge protection structure guards the LNA from damage due to static electricity. The balun converts differential signals into single-ended signals and performs impedance transformation. The matching network, located at the front-end of the LNA, converts the impedance to 50 $\Omega$, minimizing impedance matching issues at the LNA input port.
@@
For G@A, the LNA design was focused on providing higher gain at a low power consumption. The design delivers a stable gain of about 30 dB for each of the X, Y, Z inputs, in the same range as GP300, 50–200 MHz. Like in GP300, the board is powered with 5 V by the GRAND FEB; each channel draws 40 mA and the total power consumption of the board is 200 mW. As shown in Fig. 7, right, the path lengths for the X, Y, and Z inputs in the LNA board are exactly the same, leading to equal signal time delays for all channels. This allows for the study of the time dependence of the polarization in an unbiased way, avoiding additional software corrections to the timing of the individual polarization directions. Given the weather conditions in Argentina, we implemented moisture countermeasures to prevent water damage: the nuts are sealed with an additional layer of silicone, and the LNA boards are coated with transparent lacquer.
@@
The main differences in the output signals provided by the two designs of the LNA board are the gain and the matching network. For GP300, the LNA has uniform matching across the full frequency range. For the G@A model, due to the high rate of anthropogenic Radio Frequency Interferences (RFIs) in the low frequency range, its matching is optimized for higher frequencies, i.e., above 80 MHz. With this design choice, we avoid saturation of the LNA by reducing the gain at the low frequency end while keeping the relative phase delay small. The difference in maximum gain for the LNA boards is compensated for by the amplifier in the GRAND FEB, as described in the next section.
3.3 Front-End Board
@@
Figure 8: The GRAND Front-End Board (FEB). The main components: (1) signal input and filter chain, with four inputs available; (2) analog-to-digital (ADC) chip; (3) system on chip: field-programmable gate array (FPGA) and central processing unit (CPU) ; (4) Global Positioning System (GPS) chip and connector; (5) Ethernet chip; (6) clock; (7) power supply for the digital part of the board; (8) power supply for the analog part of the board; and (9) power connector and switch.
@@
Figure 9: Measured response of the analog filtering on the prototype GRAND FEB with a set 4 dB VGA gain. The plateau at around 4 dB indicates that the signal is properly amplified within the GRAND working frequency range of 50–200 MHz.
@@
Figure 8 shows the GRAND FEB prototype used in GP300 and G@A. We describe its primary components below.
Analog chain.—The analog chain, component (1) in Fig. 8, combines several functionalities. The analog signals, coming from the LNA, are first shaped through a fifth-order elliptical 30–200 MHz-bandpass filter. Then, they pass through the Variable Gain Amplifier (VGA), whose setting is controlled by the digital part of the board. The gain can vary between -10 and +20 dB. Figure 9 shows the measured analog response of the FEB board for a set VGA gain of 4 dB. In the GRAND working frequency range of 50–200 MHz, a plateau is observed at the expected 4 dB, with frequency dependent losses of at most half a dB. In GP300, we use a VGA gain of 20 dB; in G@A, we use a 0 dB gain. This choice compensates for the differences in the LNA gain and ensures that the final signal output has a similar total gain for both prototype arrays. The FEB has four analog chain signal inputs, one for each antenna polarization, and a spare that typically remains unused. The analog chain also supplies power to the front-end LNA boards through a bias-T.
@@
ADC chip.—The FEB digitizes the signals after they pass the analog chain. The digitization uses a 14-bit 500 MSPS, 4-channel Analog-to-Digital Converter (ADC; Analog Devices, AD9694), component (2) in Fig. 8, in differential mode with an input voltage range of $\pm 0.9$ V. This setting allows us to measure the background noise at around 30 ADC counts, leaving a dynamic range of about 9 effective bits. This yields a compromise between accurate noise determination and maximum signal detection that allows us to use the variations in the Galactic noise as a monitoring and calibration tool. The 500 MHz clock used by the ADC is generated locally on the board. It is not synchronized between different DUs.
@@
FPGA and CPU chip.—The digitized signals are moved into a System on a Chip (SoC) (Xilinx, Zynq Ultrascale+, XCZU7CG-2FBVB900E), component (3) in Fig. 8. It consists of a Field-Programmable Gate Array (FPGA) combined with four embedded hardcore Central Processing Units (CPUs), two ARM-A53 and two ARM-R5. The SoC performs tasks such as event triggering, event building, communication, and buffering. These functions are divided into tasks performed by the firmware that runs on the FPGA, and tasks performed by the Local Data Acquisition (DAQ) software running in the Petalinux (embedded Linux system) in the CPU—described in Section 4. Double Data Rate (DDR) memory, with a total size of 512 MB, is available to be used by the FPGA and CPU chip.
GPS and data transfer.— Timing and temperature information is obtained from the Trimble RES SMT360 GPS chip, component (4) in Fig.8, which has an accuracy of 15 ns. Wireless data transfer between the Detection Unit (DU) and the Data Acquisition (DAQ) computer is performed through the Ethernet chip, component (5) in Fig. 8.
@@
The FEB box.—As described in Section 3.1, the FEB is stored in a sealed metal box. The FEB box also acts as a Faraday cage, preventing the electronic noise generated by the FEB components from interfering with the measurements. Due to the high temperatures during the summer at the GP300 and G@A sites, it was necessary to adapt the FEB boxes to prevent the board from overheating. In GP300, a heat dissipation structure was added to the lid of the box. In G@A, the top lids of the FEB boxes were replaced with a metal mesh. A mesh could not be used in GP300 because sandstorms in the Gobi desert would introduce sand into the electronics. In addition, G@A was fitted with a firmware controlled automatic hibernation based on the ADC chip temperature.
3.4 Overall hardware system performance